OpenStudy (anonymous):

Find the x and y intercepts. f(x)=4^2+20x+25 Y= x=

OpenStudy (campbell_st):

the y intercept occurs when x = 0 so substitute x = 0... and find y x intercepts occur when y = 0 to let f(x) = 0 and solve for x is your equation \[f(x) = 4x^2 + 20x + 25 \] if it is then you have \[4x^2 + 20x + 25 = 0\] this is a perfect square as 4x^2 = (2x)^2 and 25 = (5)^2 so it can be factorised to \[(x + 5)^2 = 0\] just solve for x to get the x-intercept

OpenStudy (anonymous):

x=-5?

OpenStudy (campbell_st):

yes thats the x-intercept

OpenStudy (anonymous):

How do I get y?

OpenStudy (campbell_st):

let x = 0 so y = 4(0)^2 + 20(0) + 25

OpenStudy (campbell_st):

your y intercept is y = 25

OpenStudy (anonymous):

Is y 25??? I'm confused

OpenStudy (anonymous):

@campbell_st I think you got the x- intercept wrong 4x^2 + 20x + 25 = 0 (divide by 4) x^2 + 5x + 25/4 = 0 (group the x's) x^2 +5x = -25/4 (half x-term coefficient and square) 5 = 5/2 = 25/4 (add to both sides) x^2 + 5x + 25/4 = -25/4 + 25/4 x^2 + 5x + 25/4 = 0 \[\left( x+\frac{5}{2} \right)^2 = 0\] (square both sides and solve for x) \[\Large x = -\frac{5}{2} \pm \sqrt{0} \longrightarrow x _{1,2} = -\frac{5}{2}\] So the x-intercept should be -5/2

OpenStudy (campbell_st):

oops (2x +5)^2 =0 x = -5/2 and y = 25

OpenStudy (anonymous):

Thanks!!! ;-)